OpenStudy (anonymous):

Solve for x. 9(x - 2) = 18

OpenStudy (anonymous):

First, you'll pass the nine to the right hand side. Whatever it is doing there, it will be doing the opposite on the other side. So in this case: 9(x-2) = 18, 9 is multiplying. So you'd get: x-2= 18/9. Then for the two, it is subtracting so passing it along would make it: X= 18/9 + 2. I'll leave the rest to you.

OpenStudy (anonymous):

but I need help

OpenStudy (anonymous):

What is your question?

OpenStudy (anonymous):

Just calculate 18/9 + 2 and you will have solved for X.

OpenStudy (anonymous):

is it4 ?

OpenStudy (anonymous):

Yes. X = 4.

OpenStudy (anonymous):

If you have trouble with fractions, it is very easy work. Tutorials are online and you could ask your teacher, it would only take a minute, or use a calculator.